Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#279 closed flight processing (fixed)

GB09/05, flight day 278c, Leighton moss

Reported by: mark1 Owned by: knpa
Priority: alpha 4 low Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB09_05-2009_278c_Leighton_moss

Data arrived from ARSF via SATA disk LETTER L on 7th October 2009.

Scientific objective: /users/rsg/arsf/arsf_data/2009/ARSF_Applications/GB09-05.pdf

Priority: alpha-4-low

PI: A. Blackburn

Sensors:

  • Eagle (09/03/10)
  • Hawk (09/03/10)
  • Leica LIDAR (28/04/10) (see ticket #270)

Change History (23)

comment:1 Changed 10 years ago by mark1

  • Description modified (diff)

This project has been split from GB09/05 278b (ticket #270) and contains duplicates of the applanix and rinex data but separate Eagle/Hawk and logsheets. The Leica Lidar data has remained in GB09/05 278b since its difficult to tell where it should be split or which project areas it relates to.

comment:2 Changed 10 years ago by mark1

  • Description modified (diff)

comment:3 Changed 10 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-2.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-3.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-4.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-5.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-6.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-7.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-8.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278c_Leighton_moss/hawk/SWIR278c-09-9.hdr: predicted length half recorded length, frame rate doubling problem?

comment:4 Changed 10 years ago by chrfi

  • Owner set to chrfi
  • Status changed from new to accepted

comment:5 Changed 10 years ago by chrfi

awaiting vectors

comment:6 Changed 10 years ago by chrfi

users dosen't want to wait for vectors.

sct offsets,

flightline eagle hawk
1 0.020.88
2 -0.070.60
3 0.000.60
4 0.000.15
5 0.000.52
6 0.060.55
7 0.000.47
8 0.040.52
9 -0.040.23

comment:7 Changed 10 years ago by mark1

vectors arrived 02/03/2010

comment:8 Changed 10 years ago by mark1

Delivery checked Hyperspectral delivery. Viewed all data and tested example commands, looks good. Few edits made to read_me file. Need to change flight line numbers - then ready to deliver.

comment:9 Changed 10 years ago by chrfi

eagle and hawk data dispatched, email sent, hard drive table updated

comment:10 Changed 10 years ago by chrfi

forgot to mention, dispatched 9th March

comment:11 Changed 9 years ago by mark1

LIDAR delivery has been checked. Minor corrections made to read me and filenames. Needs comparison to vectors and probably best to split the DEM and screenshots into 2 separate ones rather than one big one.

comment:12 Changed 9 years ago by chrfi

lidar roll offsets

154230 -0.403
154612 -0.0393
155020 -0.00390
155413 -0.00393
155802 -0.00373
160344 -0.00344
160717 -0.00333
161214 -0.00359
161543 -0.00354

comment:13 Changed 9 years ago by chrfi

lidar dispatched 28/04/2010

comment:14 Changed 9 years ago by chrfi

  • Resolution set to fixed
  • Status changed from accepted to closed

comment:15 Changed 9 years ago by chrfi

  • Resolution fixed deleted
  • Status changed from closed to reopened

my mistake, still needs rsyncing and archiving.
waiting for space on bowhead

comment:16 Changed 9 years ago by knpa

  • Description modified (diff)
  • Owner changed from chrfi to knpa
  • Status changed from reopened to assigned

Begun archiving.

comment:17 Changed 9 years ago by knpa

No webcam data for this project.

comment:18 Changed 9 years ago by knpa

Hawk runscripts missing.

comment:19 Changed 9 years ago by knpa

Raw hawk unusually has one nav file that covers all the lines.

comment:20 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving

Archived and backed up by NEODC.

Moved to archive disk.

comment:21 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:22 Changed 3 years ago by lah

Tidying

There is no reprocessed hyperspectral delivery, but there are reprocessed files, so I've created a new delivery out of these and added a readme. This will need a delivery check before archiving. Oddly there are no fodis data. I have not deleted any of the hyperspectral processing files yet - will wait until after the delivery is checked.

comment:23 Changed 2 years ago by asm

Archiving

Archived and it is available from CEDA. Re-processed hyperspectral data at ​http://browse.ceda.ac.uk/browse/neodc/arsf/2009/GB09_05/GB09_05-2009_278c_Leighton_moss/hyperspectral-reprocessed

Plymouth 12th May 2017

Note: See TracTickets for help on using tickets.